An automobile manufacturer estimates that total variable costs will be $500 million and total fixed costs will be $1 billion in the next year. In setting price it is assumed that sales will be 80 percent of the firm’s 125,000 vehicle per year capacity, or 100,000 units. The target rate of return is 10 percent, which is to be earned on an investment of $2 billion. If prices are set on a cost-plus basis, what prices should be charged for each automobile?
1
Expert's answer
2020-11-17T07:37:18-0500

A target return of 10% OR $2 billion requires that the firm earn a total profit of $200 million. Computations are based sales of $100,000 vehicles. Thus, AFC, AVC, and profit per unit are:

AFC=$1000,000,000/100,000=$10,000

AVC=$500,000,000/100,000=$5000

AC=AFC+AVC=$10,000+$5000=$15,000

Per Unit profit=Total Profit/Total Units=$200,000,000/100,000=$2,000

Hence, if profit of $2000 per vehicle is added to the $15,000 In cost, the price will be $17000 per car.